RC meeting and Run Plan, Wednesday November 2nd
From SBS wiki
Return to the RC Meeting Calendar
RC Meeting Information
Time:at 3:00pm today
- Location: Counting house conference room
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
- Yesterday, it was proposed that we shift this meeting to 3:30 every day, rather than 3:00 with 3:30 on Monday (to avoid Hall C conflict). Any concerns/objections?
- Current Polarization: N/A
- Requesting hourly screenshots of Target Strip Charts (and keep an eye on these during shift) - Should be removed after configuration change ended
- Need one more set of measurements to be 100% confident in getting good results from first kinematic point - will take additional measurements on swing/owl shifts (target expert + TO)
- Hall went to restricted access this morning to allow work.
- HCal move is underway; update from Jessie
- More tests of the new VME crate (BPMs and more)
- Today was beam studies; accelerator setting up separator for 3-pass delivery
- Expect to ask for beam delivery on Saturday or Sunday; details depend on timing of work in the hall. Hopefully this will allow for extra target tests in the new configuration before data taking.
- Other Subsystem updates?
- Hall back in power permit tonight for magnet-on target measurements.
- Plan on more measurements in power permit overnight, as needed
- Tomorrow morning, move to Restricted Access (pushed back from yesterday)
- Radcon will come in ~7:30 to bag up contaminated items in target enclosure; prepare for work
- Target group will prepare target enclosure for field measurements and then make the measurements
- During some of this time, the BB/SBS magnets will be ON during the Controlled Access (under special test procedures). Check with RC before going down
- Hall open for other work, e.g. oven interlock box installation (which I mistakenly listed as already done yesterday)
- Any other access needed tonight, tomorrow, or Thursday?
Experimental progress (for MCC planning meeting): Based on UPDATED Run Coordinator instructions (new this week)
- 6.1% of accumulated charge for PRODUCTION data (13.55 Coulombs out of 232)
- 27% Experimental run time (38 out of 141 days)
- Long term "to do" items; not urgent, some can be done when 3He target not available
- (Luminosity scans with GEMs)
- (Opportunistic BCM calibration (higher current) before pass change?)
- (INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run) UPDATE: CAN BE DONE USING THE STAND-ALONE DAQ, SO NO LONGER ON THE OPPORTUNISTIC LIST.
- (Misc. GEM commissioning items:)https://sbs.jlab.org/wiki/index.php/CommissioningPlan#BB.2FSBS_GEM_Commissioning_.28Continuation.29)
Current Run Plan
'Need updated 'standard' run plan for 3-pass data taking ------
NMR measurement Restore beam
Run Plan 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour. 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization. 3.) Return to step 1.)
- Keep an eye out for drifting BPM positions
- Log all DAQ crashes in the excel worksheet along with the reason for the crash
- If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: , If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
- Cross check the rate with what is expected: 
- Cross check with the expected ion chamber readings: 
Notices in Effect for Shift Crew
- Power down SBS, BB, and Detector HV before controlled access and allowing RadCon in the hall
- Communicate to the TO that the SBS and BB magnets will be powered down, so they can update the holding field
- See the beam recovery procedures at: HOW_TOs#Beam_Recovery_Procedure
- See the instructions to set the raster at: HOW_TOs#How_to_ask_MCC_to_set_the_desired_raster_size_of_6_mm
- Shift Openings: Monday (morning) OWL, Tuesday OWL and SWING.